A uniform metre rule of mass 90g is pivoted at the 40cm mark.if the rule is in equilibrium with an unknown mass m place at the 10cm mark and a 72g mass at the 70cm mark , determine m
1
Expert's answer
2020-01-13T10:30:46-0500


The equilibrium condition


"Mgd=m_1gd_1+m_2gd_2"

Hence

"M\times 30=72\times 30+90\times 10"

"M=102 \:\rm g"